# State of practice: evaluating GPU performance of state vector and tensor network methods

### v1

@@ V-C Lessons Learnt
Following the hereby presented results and the statistics collected and discussed in Subsections V-A and V-B, we introduce in Figure 10 a decision diagram to be used for picking the most suitable simulation backend according to the properties of the quantum circuit. As we have observed, circuits which are characterised by an entanglement variance score greater than $0.2$ have a highly unbalanced structure, reducing the efficacy of the pathfinding algorithm for tensor network contraction. Likewise, if more than half of the circuit’s gates are double qubit gates, which amounts to a entanglement ratio score greater than $0.5$, the tensor network approach starts to scale poorly, as stated in Observation 4. This is due to the presence of large tensors early on during the contraction, slowing down the overall process. In both of the previous cases is thus suggested to use a state vector solver. However, if the first two conditions are not met, one must check for the program communication and the critical depth scores. If they are opposite to one another, with one being smaller than $0.2$ and the other being larger than $0.9$, then the best simulation method is the tensor network contraction. This is due to the fact that, if the previous conditions about the entanglement ratio and the entanglement variance scores are met, a program communication score greater than $0.9$ and a critical depth score lower than $0.2$ indicate a circuit structure in which most qubits interact with each other, but there are little to no repeated interactions. On the other hand, a program communication score smaller than $0.15$ together with a critical depth score larger than $0.9$ indicate a circuit structure composed of many chained two-qubit interactions among the same pairs of qubits. In both cases the tensor network becomes a pseudo-regular grid, which can be efficiently contracted whilst keeping intermediate tensor sizes at bay, leading to time performance gains of up to one order of magnitude.
@@ VI Conclusions
The paper characterised the performance of a selected suite of relevant quantum circuits when simulated on state of the art simulator backends and high performance hardware. At first, the circuits have been characterised according to objective metrics that could describe their topological structure, to later correlate them to the performance of the simulators. The results point towards the fact that state vector simulators become heavily communication bound as soon as the size of the statevector exceeds that of the GPU’s cache. The tensor network contraction has proven to perform better in circuits that have well distributed entanglement among the qubits, with a low overall number of two qubit gates with respect to the total number of gates.
It is reasonable to assume that by improving the memory access pattern of a state vector backend, one may improve the time performance of any benchmark to be simulated. On the other hand, the tensor network backend proved to be highly dependent on the efficacy of the pathfinding algorithm. The overall performance of this second backend can already outpace the state vector simulator in some of the benchmark problems, whilst keeping a comparable, albeit slower performance in the remainder of the test runs. It is reasonable to assume that the development of an optimised pathfinding algorithm for these harder to tackle topologies may in fact push the performance of the tensor contraction backend to become the overall fastest simulation approach for quantum circuits. Moreover, implementing a GPU scalable version of the pathfinding algorithm may provide better performance over larger tensor networks without incurring in a computation time trade off. This would indeed be favourable, as tensor network contraction does not suffer from exponential memory requirements as the state vector approach, thus enabling the validation of larger quantum circuits.

### v2

@@ V-E Lessons learnt
As we have observed, circuits which are characterised by an entanglement variance score greater than $0.2$ have a highly unbalanced structure, reducing the efficacy of the pathfinding algorithm for tensor network contraction. Likewise, if more than half of the circuit’s gates are double qubit gates, which amounts to a entanglement ratio score greater than $0.5$, the tensor network approach starts to scale poorly, as stated in Observation 4. This is due to the presence of large tensors early on during the contraction, slowing down the overall process. In both of the previous cases is thus suggested to use a state vector solver. However, if these first two conditions are not met, one must check for the program communication and the critical depth scores. If they are opposite to one another, with one being smaller than $0.2$ and the other being larger than $0.9$, then the best simulation method is the tensor network contraction. This is due to the fact that, if the previous conditions about the entanglement ratio and the entanglement variance scores are met, a program communication score greater than $0.9$ and a critical depth score lower than $0.2$ indicate a circuit structure in which most qubits interact with each other, but there are little to no repeated interactions. On the other hand, a program communication score smaller than $0.15$ together with a critical depth score larger than $0.9$ indicate a circuit structure composed of many chained two-qubit interactions among the same pairs of qubits. In both cases the tensor network becomes a pseudo-regular grid, which can be efficiently contracted whilst keeping intermediate tensor sizes at bay, leading to time performance gains of up to one order of magnitude.
Program communication is especially important in determining whether a quantum circuit contraction can be efficiently contracted in a distributed sliced setting. Those circuits display superlinear speedups with respect to the available compute resources. Moreover, we showed how is is only justifiable to provide additional pathfinding resources to unbounded quantum circuit tensor contractions, as they can provide further speedups, whilst all other circuits can save on using additional resources on pathfinding.
@@ VI Conclusions
The paper characterised the performance of a selected suite of relevant quantum circuits when simulated on state of the art simulator backends and high performance hardware. At first, the circuits have been characterised according to objective metrics that could describe their topological structure, to later correlate them to the performance of the simulators. The results point towards the fact that statevector simulators become heavily communication bound as soon as the size of the statevector exceeds that of the GPU’s cache. The tensor network contraction has proven to perform better in circuits that have well distributed entanglement among the qubits, with a low overall number of two qubit gates in relation to the total number of gates.
It is reasonable to assume that by improving the memory access pattern of a state vector backend, one may improve the time performance of any benchmark to be simulated. The tensor network backend proved to be highly dependent on the efficacy of the pathfinding algorithm. The overall performance of this second backend can already outpace the state vector simulator in some of the benchmark problems, whilst keeping a comparable, albeit slower performance in the remainder of the test runs. It is reasonable to assume that the development of further pathfinding optimisations for these harder to tackle topologies may push the performance of the tensor contraction backend to become the fastest simulation approach for quantum circuits, for example by limiting the growth of intermediate tensors during contraction. This would let us leverage the fact that the representation of tensor networks in memory enables the validation of larger quantum circuits and computers. Moreover, the promising scaling of distributed pathfinding and sliced contraction approaches over large tensor networks have the chance to further close the gap on simulating real quantum computer. We highlight the absence of a class of real-world quantum algorithms with parameterisable contraction complexity. The development of such a class of algorithms would provide means for measuring weak scaling performance of distributed tensor network contraction libraries, easing performance comparisons between quantum and classical systems. Future works could investigate the applicability of GPU accelerated pathfinding algorithms, so as to further improve the path quality and reduce the overall contraction time in unbounded problems.